Brooks Clinic is considering investing in new heart-monitoring equipment. It has two options. Option A would have an initial lower cost but would require a significant expenditure for rebuilding after 4 years. Option B would require no rebuilding expenditure, but its maintenance costs would be higher. Since the Option B machine is of initial higher quality, it is expected to have a salvage value at the end of its useful life. The following estimates were made of the cash flows. The company’s cost of capital is 6%.

Option A Option B
Initial Cost $170,000 $274,000
Annual Cash InFlows $72,200 $82,200
Annual Cash OutFlows $31,400 $26,700
Cost to Rebuild (end of 4 years) $50,100 $0
Salvage Value $0 $8,200
Estimated Useful Life 7 Years 7 Years

Compute the (1) net present value, (2) profitability index, and (3) internal rate of return for each option. (Hint: To solve for internal rate of return, experiment with alternative discount rates to arrive at a net present value of zero.) (If the net present value is negative, use either a negative sign preceding the number eg -45 or parentheses eg (45). Round answers for present value and IRR to 0 decimal places, e.g. 125 and round profitability index to 2 decimal places, e.g. 12.50. For calculation purposes, use 5 decimal places as displayed in the factor table provided.)

Net Present Value Profitability Index Internal Rate of Return
Option A $ %
Option B $ %

Which option should be accepted?

Option A or Option B should be accepted?
Answer #1

Answer:

Option A:

Net Present Value:

Year

Flow

Net cash flow

Present value

Computation

0

-170,000

-170,000

-170,000

1

72,200-31,400

40,800

38,490.57

40,800 / 1.06

2

72,200-31,400

40,800

36,311.85

40,800 / 1.06^2

3

72,200-31,400

40,800

34,256.47

40,800 / 1.06^3

4

72,200-31,400-50,100

-9,300

-7,366.47

-9300 / 1.06^4

5

72,200-31,400

40,800

30,488.13

40,800 / 1.06^5

6

72,200-31,400

40,800

28,762.39

40,800 / 1.06^6

7

72,200-31,400

40,800

27,134.33

40,800 / 1.06^7

Add up the present value column to get the Net present value,

Net Present Value = $18,077.27

Profitability Index:

Profitability Index = Present value of future cash flows / Initial Investment

Present value of future cash flows can be found out by adding present value column values except initial investment.

Present value of future cash flows = 38,490.57 + 36,311.85 + 34,256.47 - 7,366.47 + 30,488.13

                                                                         + 28,762.39 + 27,134.33

                                                                = 188077.27

Profitability index = 188077.27 / 170000

                                    = 1.11

Internal Rate of Return:

If cost of capital is 8% then, NPV = 5,594.90279

For 9% cost of capital, NPV = -147.62740

Interpolating, (9% - 8%) / (5594.90279 + 147.62740) * 5594.90279 = 0.9743

IRR = 8 + 0.9743 = 8.9743%

Option B:

Year

Flow

Net cash flow

Present value

Computation

0

-274,000

-274,000

-274,000

1

82,200-26,700

55,500

52,358.49

55,500 / 1.06

2

82,200-26,700

55,500

49,394.8

55,500 / 1.06^2

3

82,200-26,700

55,500

46,598.87

55,500 / 1.06^3

4

82,200-26,700

55,500

43,961.2

55,500 / 1.06^4

5

82,200-26,700

55,500

41,472.83

55,500 / 1.06^5

6

82,200-26,700

55,500

39,125.31

55,500 / 1.06^6

7

82,200-26,700+8,200

63,700

42,364.14

63,700 / 1.06^7

Net Present Value = $41,275.64

Profitability Index:

Profitability Index = Present value of future cash flows / Initial Investment

Profitability index = 315275.64 / 274000

                                    = 1.15

Internal Rate of Return:

If cost of capital is 10% then, NPV = 405.14095

For 11% cost of capital, NPV = -8523.50834

Interpolating, (11%-10%) / (405.14095 + 8523.50834) * 405.14095 = 0.04538

IRR = 10 + 0.04538 = 10.04538%

Net Present Value

Profitability Index

Internal Rate of Return

Option A

$18,077

1.11

8%

Option B

$41,276

1.15

10%

Option B should be accepted.